### v1

# Introducing sgboost: A Practical Guide and Implementation of sparse-group boosting in R

Abstract
This paper introduces the sgboost package in R, which implements sparse-group boosting for modeling high-dimensional data with natural groupings in covariates. Sparse-group boosting offers a flexible approach for both group and individual variable selection, reducing overfitting and enhancing model interpretability. The package uses regularization techniques based on the degrees of freedom of individual and group base-learners, and is designed to be used in conjunction with the mboost package. Through comparisons with existing methods and demonstration of its unique functionalities, this paper provides a practical guide on utilizing sparse-group boosting in R, accompanied by code examples to facilitate its application in various research domains. Overall, this paper serves as a valuable resource for researchers and practitioners seeking to use sparse-group boosting for efficient and interpretable high-dimensional data analysis.
keywords:
group selection; variable selection; R package; Ridge regression; within-group sparsity
articletype: Application Note
1 Introduction
Regularized regression is used to model high-dimensional data to reduce the risk of overfitting and to perform variable selection. In many cases, covariates have natural groupings, such as with gene data or categorical data often found in survey data. In such cases, one may want to select whole groups of variables or just individual parts (univariate regression model). ‘sgboost‘ implements the sparse-group boosting in R and other useful functions for sparse group model interpretation unique to boosting and visualization for group and individual variable selection. The package is available on CRAN [8]. A formula object defining group base-learners and individual base-learners is utilized for model fitting. Regularization is based on the degrees of freedom $(df(\cdot))$ of individual and group base-learners as defined in [10]. For an individual base-learner, the Regularization parameter $\lambda$, and for a group base-learners the regularization parameter $\lambda^{(g)}$ is defined, such that $df(\lambda)=\alpha$ and $df(\lambda^{(g)})=1-\alpha$. The package should be used with the ’mboost’ package [4]. Sparse-group boosting is an alternative method to sparse-group lasso, employing boosted Ridge regression [14]. Although there are many methods of variable selection, most focus on group selection only, e.g. [7], [5] and [16], or individual variable selection e.g. [1], [15] and [2]. However, it should be noted, that with group variable selection with overlapping groups, one could also end up with sparse-group variable selection. There are not many R packages implementing sparse-group variable selection. There is ’SGL’ [13] implementing [14] and ’sparsegl’ [6] with a faster implementation of the sparse-group lasso.
The goal of this paper is to provide a practical guide including the code on how to use sparse-group boosting in R and get the most out of the method. The code is presented within this manuscript and can be found also on GitHub together with the used dataset (https://github.com/FabianObster/sgboost-introduction).
2 Sparse-group boosting workflow with sample data
We first simulate the sample data and corresponding group structure with 40 equal-sized groups. Based on a linear regression model we simulate the response variable y as part of the data.frame with $n=100$ observations and $p=200$ predictor variables (each group is formed by 5 predictors).
beta <- c(
rep(5, 5), c(5, -5, 2, 0, 0), rep(-5, 5),
c(2, -3, 8, 0, 0), rep(0, (200 - 20))
X <- matrix(data = rnorm(20000, mean = 0, sd = 1), 100, 200)
df <- data.frame(X) %>%
mutate(y = X %*beta+rnorm(100, mean = 0, sd = 1)) %>%
mutate_all(function(x){as.numeric(scale(x))})
group_df <- data.frame(
group_name = rep(1:40, each = 5),
variable_name = colnames(df)[1:200]
2.1 Defining the model
Now we use the group structure to describe the sparse group boosting formula with the function create_formula(). We only need the data.frame() describing the group structure. It should contain two variables one indicating the name of the variable in the modeling data (var_name), and one indicating the group it belongs to (group_name). Additionally, we need to pass the mixing parameter alpha and the name of the outcome variable.
sgb_formula <- create_formula(
alpha = 0.4, group_df = group_df, outcome_name = "y",
group_name = "group_name", var_name = "variable_name")
This function returns an R-formula consisting of $p$ model terms defining the individual base-learners and $G$ group base-learners.
labels(terms(sgb_formula))[[1]]
## bols(X1, df = 0.4, intercept = FALSE)
labels(terms(sgb_formula))[[201]]
## bols(X1, X2, X3, X4, X5, df = 0.6, intercept = FALSE)
2.2 Fitting and tuning the model
sgboost is to be used in conjunction with the mboost package which provides many useful functions and methods which can also be used for sparse-group boosting models.
Now we pass the formula to mboost() and use the arguments as seems appropriate. The main hyperparameters are nu and mstop. For model tuning, the function cvrisk can be used and plotted. Running the cross-validation/bootstrap in parallel can speed up the process.
sgb_model <- mboost(
formula = sgb_formula, data = df,
control = boost_control(nu = 1, mstop = 600)
cv_sgb_model <- cvrisk(sgb_model)
mstop(cv_sgb_model)
## 204
plot(cv_sgb_model
Figure 1: Out of sample error depending on the boosting iteration
In this example, the lowest out-of-sample risk is obtained at 204 boosting iterations, so we only use the first 204 updates for the final model.
2.3 Interpreting and plotting a sparse-group boosting model
sgboost has useful functions to understand sparse-group boosting models and reflects that final model estimates of a specific variable in the dataset can be attributed to group base-learners as well as individual base-learners depending on the boosting iteration.
2.3.1 Variable importance
A good starting point for understanding a sparse-group boosting model is the variable importance. In the context of boosting, the variable importance can be defined as the relative contribution of each predictor to the overall reduction of the loss function (negative log-likelihood). get_varimp() returns the variable importance of each base-learner/predictor selected throughout the boosting process. In the case of the selection of an individual variable - call it $x_{1}$ as well as the group it belongs to -$x_{1},x_{2},...x_{p}$, both base-learners (predictors) will have an associated variable importance as defined before. This allows us to differentiate between the individual contribution of $x_{1}$ as its own variable and the contribution of the group $x_{1}$ belongs to. It is impossible to compute the aggregated variable importance of $x_{1}$ as it is unclear how much $x_{1}$ contributes to the group. However, the aggregated coefficients can be computed using get_coef(), which also returns the aggregated importance of all groups vs. all individual variables in a separate data.frame. With plot_varimp() one can visualize the variable importance as a barplot. Since group sizes can be large, the function allows to cut of the name of a predictor after max_char_length characters. One can indicate the maximum number of predictors to be printed through n_predictors or through the minimal variable importance a predictor has to have through prop. Through both parameters, the number of printed entries can be reduced. Note, that in this case, the relative importance of groups in the legend is based only on the plotted variables and not the ones removed. Adding information about the direction of effect sizes, one could add arrows behind the bars [9]. For groups, one can use the aggregated coefficients from get_coef().
slice(get_varimp(sgb_model =sgb_model_linear)\$varimp,1:5)
# A tibble: 5  6
  reduction blearner        predictor   selfreq type  relative_
                                                      importance
      <dbl> <chr>           <chr>       <dbl>   <chr>     <dbl>
1    0.297  bols(X1, X2,... X1, X2, ... 0.206   group     0.301
2    0.288  bols(X18, in... X18         0.0196  indi...   0.292
3    0.230  bols(X11, X1... X11, X12... 0.25    group     0.233
4    0.0414 bols(X7, int... X7          0.0784  indi...   0.0419
5    0.0392 bols(X6, int... X6          0.0833  indi...   0.0397
get_varimp(sgb_model = sgb_model_linear)\$group_importance
    # A tibble: 2  2
  type       importance
  <chr>           <dbl>
1 group           0.534
2 individual      0.466
plot_varimp(sgb_model = sgb_model_linear, n_predictors = 15)
Figure 2: Variable importance of the sparse-group boosting model for simulated data
In this example, we see that both individual variables and groups were selected and contributed to the reduction of the loss function. The most important predictor is the first group, followed by variable 18, and then followed by group three. This is in line with what was simulated, as variable 18 has the biggest beta value, and groups one and three are full groups, meaning all variables within the groups have a non-zero beta coefficient. Groups two and four have within-group sparsity, therefore they were selected as individual variables rather than groups.
2.3.2 Model coefficients
The resulting coefficients can be retrieved through get_coef() In sparse-group boosting models, a variable in a dataset can be selected as an individual variable or through a group. Therefore, there can be two associated effect sizes for the same variable. This function aggregates both and returns them in a data.frame sorted by the effect size ’effect’.
slice(get_coef(sgb_model = sgb_model)\$raw,1:5)
# A tibble: 5  5
  variable effect blearner         predictor               type
  <chr>     <dbl> <chr>            <chr>                   <chr>
1 X18       0.364 bols(X18, int... X18                     individual
2 X5        0.250 bols(X1, X2, ... X1, X2, X3, X4, X5      group
3 X15      -0.249 bols(X11, X12... X11, X12, X13, X14, X15 group
4 X4        0.234 bols(X1, X2, ... X1, X2, X3, X4, X5      group
5 X11      -0.228 bols(X11, X12... X11, X12, X13, X14, X15 group
slice(get_coef(sgb_model = sgb_model)\$aggregate,1:5)
# A tibble: 5  4
 variable effect learner               predictor
  <chr>    <dbl> <chr>                 <chr>
1 X18      0.364 bols(X18, inte...     X18
2 X15     -0.272 bols(X11, X12,...;    X11, X12, X13, X14, X15; X15
                 bols(X15, inte...
3 X5       0.250 bols(X1, X2,...       X1, X2, X3, X4, X5
4 X4       0.234 bols(X1, X2,...       X1, X2, X3, X4, X5
5 X13     -0.230 bols(X11, X12,...;    X11, X12, X13, X14, X15; X13
                 bols(X13, inte...
We see that the effect sizes differ between the two perspectives. The variable X15 for example has a more extreme model coefficient of -0.272 in the aggregated case compared to the coefficient of -0.249 derived only from the group base-learner. Consequently, also the ordering differs. X11 has a greater absolute model coefficient from the group than X13, but in the aggregated version the absolute model coefficient of X13 exceeds the one of X11.
2.3.3 Plotting model coefficients and importance
With plot_effects() we can plot the effect sizes of the sparse-group boosting model in relation to the relative importance to get an overall picture of the model. Through the parameter ’plot_type’ one can choose the type of visualization. ’radar’ refers to a radar plot using polar coordinates. Here the angle is relative to the cumulative relative importance of predictors and the radius is proportional to the effect size. ’clock’ does the same as ’radar’ but uses clock coordinates instead of polar coordinates. ’scatter’ uses the effect size as the y-coordinate and the cumulative relative importance as the x-axis in a classical Scatter plot.
plot_effects(sgb_model = sgb_model, n_predictors = 5,
base_size = 10)
plot_effects(sgb_model = sgb_model, n_predictors = 5,
plot_type = "clock", base_size = 10)
plot_effects(sgb_model = sgb_model, n_predictors = 5,
plot_type = "scatter", base_size = 10)
Figure 3: Effect size and
2.3.4 Coefficient path
plot_path calls get_coef_path() to retrieve the aggregated coefficients from a mboost object for each boosting iteration and plots it, indicating if a coefficient was updated by an individual variable or group.
plot_path(sgb_model = sgb_model)
Figure 4: Coefficient path of a sparse-group boosting model with simulated data
In the coefficient path shown in Figure 4, we see the change in model coefficients. Since the path shows the aggregated model coefficients, the path of one variable in the dataset may have both colors. This is the case with variable X1 which was first updated through the group and then also as an individual variable or with variable X15 in reverse order.
3 Real data
In this section, we will fit a sparse-group boosting model with sgboost to a real dataset. We will use behavioral ecological data and an associated group structure [11] to explain whether farmers in Chile and Tunisia are planning to take adaptive measures against climate change in the following years. We will use a logistic regression model for this binary decision. The data consists of 14 groups and 84 variables for the 801 farmers. Groups include vulnerability to climate change [12], social, biophysical, and economic assets, as well as perceptions of the farmers. After loading the data and group structure, we create the formula with mixing parameter $\alpha=0.3$. Then, we pass the formula to mboost() with 1000 boosting iterations and a learning rate of 0.3.
model_df <- readRDS(’model_df.RDS’) %>%
mutate_at(index_df\$col_names, factor)
index_df <- readRDS(’index_df.RDS’)
sgb_formula <- create_formula(
group_df = index_df, var_name = col_names’,
group_name = index’, outcome_name = S5.4’
model <- mboost(
sgb_formula, data = model_df,
family = Binomial(link = logit’),
control = boost_control(mstop = 1000, nu = 0.3)
cv_model <- cvrisk(model)
model <- model[mstop(cv_model)]
The model is stopped early after 466 Boosting iterations. We examine the coefficient path and see that in the early stage, individual base-learners were dominantly selected like the variable ’S1.8b or ’S8.11 river’ which indicates whether river irrigation is used. Many of the variables were first included as individual variables and later also through group base-learners like ’S8.1b’ or ’S2.5c proximity’ (Proximity to extreme weather events), which we also saw in the simulated data.
plot_path(model)
Figure 5: Coefficient path using the ecological dataset
In figure 6, we look at the variable importance with the default values, plotting all 27 selected predictors of which 8 are groups, the latter having a relative variable importance of 22 percent. The most important base-learner is the individual variable ’S1.8b’ indicating whether farming journals are being used and the most important group is the social asset group followed by the group consisting of the four considered regions.
plot_varimp(model)
Figure 6: Variable importance using the ecological dataset
Plotting the effect sizes of all predictors having a relative importance of greater than 1.5 percent shows the tendency for more important variables to have greater absolute effect sizes. For readability, we set the number of printed characters per variable to 6 and use the ’scatter’ version of the plot.
plot_effects(
model, plot_type = scatter’,
prop = 0.015, max_char_length = 6
Figure 7: Coefficient plot using the ecological dataset
4 The sparse-group boosting algorithm
General functional gradient descent Algorithm ([3])
  1. 1.
    Define base-learners of the structure $h:\mathbb{R}^{n\times p}\to\mathbb{R}$
  2. 2.
    Initialize $m=0$ and $\widehat{f}^{[0]}\equiv 0$ or $\widehat{f}^{[0]}\equiv\overline{y}$
  3. 3.
    Set $m=m+1$ and compute the negative gradient $\frac{\partial}{\partial f}l(y,f)$ and evaluate it at $\widehat{f}^{[m-1]}$. Doing this yields the pseudo-residuals $u_{1},...,u_{n}$ with
    $u_{i}^{[m]}=\frac{\partial}{\partial f}l(y_{i},f)|_{f=\widehat{f}^{[m-1]}},$
    for all $i=1,...,n$
  4. 4.
    Fit the base-learner $h$ with the response $(u_{1}^{[m]},...,u_{n}^{[m]})$ to the data. This yields $\widehat{h}^{[m]}$, which is an approximation of the negative gradient
  5. 5.
    Update
    $\widehat{f}^{[m]}=\widehat{f}^{[m-1]}+\eta\cdot\widehat{h}^{[m]}$
    here $\eta$ can be seen as learning rate with $\eta\in]0,1[$
  6. 6.
    Repeat steps 3,4 and 5 until $m=M$
‘sgboost‘ uses the following adjustment to the general functional gradient descent algorithm. The sparse-group-boosting optimization problem is formulated as the sum of mean squared error (linear regression) or negative log-likelihood (generalized linear models e.g. logistic regression) within each boosting iteration in steps 3 and 4. The base-learner with the smallest evaluated loss function will be selected.
Linear regression:
  • Group base-learner:
    $\operatorname*{arg\,min}_{g\leq G}\min_{\beta^{(g)}\in\mathbb{R}^{p}}\left(\frac{1}{2n}\rVert u^{[m]}-X^{(g)}\beta^{(g)}\rVert_{2}^{2}+\lambda^{(g)}\rVert\beta^{(g)}\rVert_{2}^{2}\right).$
  • Individual base-learner:
    $\operatorname*{arg\,min}_{g\leq G,j\leq p_{g}}\min_{\beta^{(g)}_{j}\in\mathbb{R}}\left(\frac{1}{2n}\rVert u^{[m]}-X^{(g)}_{j}\beta^{(g)}_{j}\rVert_{2}^{2}+\lambda^{(g)}_{j}\rVert\beta^{(g)}_{j}\rVert_{2}^{2}\right).$
If the group base-learner with the smallest loss as defined above has a smaller loss than the individual base-learner with the smallest loss then the group base-learner will be selected in the update and vice versa.
Generalized linear models:
  • Group base-learners:
    $\operatorname*{arg\,min}_{g\in G}\min_{\beta^{(g)}\in\mathbb{R}}\left(-\text{loglik}(u^{[m]},X^{(g)},\beta^{(g)})+\lambda^{(g)}\rVert\beta^{(g)}\rVert_{2}^{2}\right).$
  • Individual base-learners:
    $\operatorname*{arg\,min}_{g\in G,j\in p_{g}}\min_{\beta^{(g)}_{j}\in\mathbb{R}}\left(-\text{loglik}(u^{[m]},X^{(g)}_{j},\beta^{(g)}_{j})+\lambda^{(g)}_{j}\rVert\beta^{(g)}_{j}\rVert_{2}^{2}\right).$
where
  • $df(\lambda^{(g)}_{j})=\alpha$, $df(\lambda^{(g)})=1-\alpha$
  • $df(\lambda)=tr(2H_{\lambda}-H_{\lambda}^{T}H_{\lambda})$ and $H_{\lambda}$ is the Ridge Hat matrix of a base-learner
  • $X^{(g)}$ is the submatrix of $X$ with columns corresponding to group $g$ and $X^{(g)}_{j}$ the $j$-th column of group $g$.
  • $\beta^{(g)}$ and $\beta^{(g)}_{j}$ are the corresponding parameter vectors of the features in group $g$.
  • $\alpha$ adjusts the weight between group and individual base-learners.
  • $\lambda^{(g)},\lambda^{(g)}_{j}$ Ridge penalty parameters.
  • $u^{[m]}$ is the current negative gradient vector from the previous boosting iteration.
Funding
This research is funded by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr. dtec.bw is funded by the European Union – NextGenerationEU. All statements expressed in this article are the authors’ and do not reflect the official opinions or policies of the authors’ host affiliations or any of the supporting institutions.